For the [TeCl6]^3- and [Tc(NCS)6]^3-, the crystal field splitting energies are 16,500 cm^-1
and 31,000 cm^-1, respectively, while the electron-pairing energy (P) for the Tc^+3 ion is
21,000 cm^1-.

a. Write the d electronic structures of the two complexes:

[TeCl6]^3-

[Tc(NCS)6]^3-

b. Calculate the CFSE for these two complexes (in cm):

[TeCl6]^3-

[Tc(NCS)6]^3-

Show CFSE calculations.
Answer #1

a. The d-electronic configuration of [TcCl6]3- = t2g3eg1

Explanation: The chloride ligand (Cl-) is a weak field ligand, as a result, the crystal field splitting is less, so the fourth electron occupies eg level after the t2g level is half-filled.

The d-electronic configuration of [Tc(NCS)6]3- = t2g4eg0

Explanation: The thiocyanate ligand (NCS-) is a strong field ligand, as a result, the crystal field splitting is high, so the fourth electron pairs up in the t2g level itself, after the t2g level is half-filled.

b. CFSE for [TcCl6]3- = 3*(-0.4 o) + 1*(0.6 o) = -0.6 o = -0.6*16500 cm-1 = -9900 cm-1

CFSE for [Tc(NCS)6]3- = 4*(-0.4 o) + 1*(P) = -1.6 o + P = -1.6*31000 cm-1 + 21000 cm-1 = -28600 cm-1